# Unsourced Random Access: A Comprehensive Survey

### v2

@@ V URA over Fading Channels
Since GMAC is not a very realistic channel model for wireless communications, most works in the URA literature consider fading channels, the majority of them dedicated to Rayleigh fading channels. In this section, we first review the bounds on the performance of fading URA solutions that have been provided in the literature. Then, we cover the URA solutions that present low-complexity coding schemes to achieve those bounds. To this end, in order to explain the fading URA approaches in a simpler setup for better understanding, we first cover the single-antenna fading solutions, then, we turn our attention to the MIMO fading ones.
@@ V-A Performance Bounds
In [85, 83, 84, 80], the authors derive some performance bounds for the URA setup over a Rayleigh fading channel. In one of the main analytical results, the authors of [80] present a converse bound for fading URA based on the results in [81] and the meta-converse bound in [82], as stated below.
Theorem 2 (Converse bound [80]).
Let
$L_{n}=n\hskip 2.84526pt\log(1+PG)+\sum\limits_{i=1}^{n}\left(1-\absolutevalue{\sqrt{PG}Z_{i}-\sqrt{1+PG}}^{2}\right),$ (14)
$S_{n}=n\hskip 2.84526pt\log(1+PG)\sum\limits_{i=1}^{n}\left(1-\frac{\absolutevalue{\sqrt{PG}Z_{i}-1}^{2}}{1+PG}\right),$ (15)
where $G=\norm{h}^{2}$ and $Z_{i}\sim\mathcal{CN}(0,1)$. Then, for every $n$ and $10<\epsilon<$ 1, any $(M,n-1,\epsilon)$ code for the quasi-static $K_{a}$ MA satisfies
$\log(M)\leq\log(K_{a})+\log\frac{1}{\mathbb{P}(L_{n}\geq n\gamma_{n})},$ (16)
where $\gamma_{n}$ is the solution of $\mathbb{P}\left[S_{n}\leq n\gamma_{n}\right]=\epsilon$.
For the proof, first, consider the fact that the converse bound when channel state information (CSI) is available at the receiver is a converse for the case with no CSI case as well. Moreover, by symmetry, it is enough to lower bound the probability that a particular user’s message is not in the decoded list. Finally, assuming that the decoder knows the codewords of all the other users (which makes this in essence a single-user bound), the rest of the proof becomes quite straightforward, which can be found in [80]. The authors numerically show that the $E_{b}/N_{0}$ requirements for a fading channel are necessarily higher compared to the AWGN channel. They also mention that, for the common system parameters in a fading URA setup, a Fano-type converse bound is worse than the one presented in Theorem 2.
In [83] and [84], achievability and converse bounds are proposed for the cases of known and unknown CSI in an asymptotic setting, where the number of users grows linearly with the blocklength. The authors in [80] provide both asymptotic and non-asymptotic achievability and converse bounds for the case of unknown CSI. In [85], authors derive an achievability bound for the URA scheme over the Rayleigh fading channel by considering slotted structure and perfect SIC in their analysis, which is tighter than the other existing bounds. However, since the channel coefficients used in a SIC process are estimated, hence imperfect, the assumption of perfect SIC is not realistic, which makes the result in [85] an approximation result rather than an achievability one.
Moreover, in [86, 87], achievability and converse bounds are obtained for URA over MIMO Rayleigh fading channels, whose ML-based achievability bound is obtained by employing a projection decoder presented in [88].
@@ V-B Single-Antenna Fading Channels
We first take a look at the existing URA fading solutions that assume a single-antenna receiver. The works in [72, 94, 96, 85, 80, 89, 90, 92, 95, 93, 91] introduce practical coding schemes tailored for the URA scheme on single-antenna fading channels. Within this collection of works, some assume that the received signals from all users are synchronized [72, 85, 80, 89], while others take into account asynchronous transmission [95, 90, 92, 93, 91]. To estimate the delay of asynchronous signals, [95, 90, 92, 91] use OFDM transmission to convert the time delay to the phase shift in the frequency domain, while the authors in [93] consider time-domain transmission where the delay of a signal is detected via correlation-based approach.
Focusing on the Rayleigh fading channels, in the following, we present the existing fading URA solutions, categorized based on their general approach.
Figure 10: Another representation of encoding (top) and decoding (bottom) structures of a CCS scheme with 4 slots [60].
@@ V-B1 CCS-based schemes
As explained in the previous section, CCS was originally proposed for URA over GMAC. Nevertheless, it has been applied to the fading scenario as well in several works [89, 93, 96]. In Fig. 10, another representation of the encoding (upper half of the figure) and decoding (lower half) structures of a CCS-based scheme with four slots are depicted, where each user’s signal is divided into four parts, with each one being appended to a parity-check bit, which is then mapped to a codebook to construct the pilot of the user in a slot. Also, at the receiver, sub-messages of different users are detected at each slot, and corresponding ones in different slots are stitched together to form the complete message sequence of a user. In [89], the authors modify the tree decoder of the GMAC CCS scheme in [60] by replacing it with two codes capable of correcting $t$ errors, in which the first one modifies the conventional tree code by giving it the ability to recover $t$ errors (called $t$-tree), and the second one is a Reed-Solomon (RS) code. In [96], the outer encoded message bits are partitioned into sub-blocks, and each sub-block is transmitted in a different slot after CS encoding, with AMP at the receiver. Moreover, in [93], the authors utilize CCS for the asynchronous URA model. More specifically, during the encoding stage, they append $T_{z}$ additional zeros to the end of each CS sequence, where $T_{z}$ represents the maximum possible delay of the system. At the receiving end, the decoder takes into account all potential delays, ranging from one symbol to a maximum of $T_{z}$-symbol delays when performing the CS decoding.
@@ V-B2 Joint decoding-based schemes
In [80], a joint decoding approach is employed. For encoding, this scheme partitions the time frame into distinct slots. Each user transmits its LDPC codeword through a randomly chosen slot. As previously mentioned, slot allocation offers two significant benefits: it reduces the computational load on the decoder by reducing the length of the LDPC signal, and mitigates per-slot interference by distributing users across slots. This scheme utilizes an iterative belief propagation (BP) decoder, as illustrated in Fig. 11. The graph incorporates four distinct node types: variable nodes (in red), and check nodes (in blue), which constitute the traditional LDPC decoder; functional nodes (in green) corresponding to the symbols of the received signal; and a fourth type (in magenta) corresponding to the fading coefficients. This graph allows them to simultaneously update the estimated values of both the channel coefficients and the symbols for each user.
Figure 11: Iterative BP decoder for a two-users LDPC-based scheme with joint decoding [80].
@@ V-B3 TIN-SIC-based schemes
A lower-complexity alternative to joint decoding is TIN-based decoding. Similar to GMAC, the TIN solution for fading channels operates under the assumption of the presence of just a single user and treats other interfering users as noise. Consequently, this approach results in an exceptionally low average SNR due to the fact that it encounters interference-plus-noise as the system’s overall noise level, leading to a very low performance. However, when combined with an appropriate channel code and SIC, it can lead to good performance.
@@
Paper Approach Description Synchron. vs. Asynch.
@@
Kowshik et al., 2019&2021 [83, 84] - Achievable and converse bounds for known and unknown CSI in an asymptotic setting + comparing with CDMA, TDMA, and FDMA. Synchron.
Andreev et al., 2019 [90, 91] TIN-SIC A new coding scheme containing LDPC code, OFDM, slotting, and SIC. Asynch.
Andreev et al., 2020 [85] TIN-SIC Non-asymptotic achievability and converse bounds for unknown CSI + a new coding scheme containing slotting, polar code, and SIC. Synchron.
@@
Kowshik et al., 2019&2020 [80] joint decoding Achievability and converse bounds for known and unknown CSI + a scheme containing LDPC code, slotting, and iterative joint decoding. Synchron.
Amalladinne et al., 2019 [93] CCS-based They extend CCS in [60] to accommodate asynchronous scenarios. Asynch.
@@
Andreev et al., 2022 [89] CCS-based Improving CCS in [60] by replacing the conventional outer tree code with 1) a code capable of correcting $t$ errors, and 2) a Reed–Solomon code. Synchron.
@@
Yun et al., 2024 [96] CCS-based A slotted scheme including CCS and AMP in a block fading model. Synchron.
@@
Nassaji et al., 2022 [94] Spreading-based A non-slotted scheme including, preamble transmission, channel coding, repetition, permutation, scrambling, and SIC. Synchron.
Chen et al., 2017 [92] OFDM+ SIC A new coding scheme containing OFDM and SIC. Asynch.
@@
Ozates et al., 2023 [95] OFDM+SIC A new coding scheme in the frequency-selective channel, containing OFDM, pilot transmission, slotting, polar code, and SIC. Asynch.
@@
TIN-SIC with LDPC coding: In [90], building upon their asynchronous schemes presented in [80], the authors present a low-complexity coding scheme which incorporates orthogonal frequency division multiplexing (OFDM) techniques to mitigate the effects of the asynchronous transmission by converting the time delays into phase shifts, and employing SIC to decode the user packets. More numerical results for this scheme is presented in [91].
@@
TIN-SIC with Polar Coding: In [85], a TIN-SIC solution in conjunction with polar coding is presented, where the time frame in this approach is divided into $S$ slots. Each user randomly selects a slot to transmit its polar coded and modulated signal. For decoding, the scheme considers a TIN strategy followed by an SIC. This is because different users in a slot are prone to different fading coefficients; some of them with larger coefficients are more likely to be decoded by TIN. After decoding the strongest user’s signal while treating others as noise, its channel coefficient is estimated using the orthogonal matching pursuit (OMP) technique. Given the estimated channel coefficient and the decoded signal of a user, its contribution can be removed from the received signal using SIC. This strategy continues until no signal is successfully detected during an iteration. It is important to emphasize that the determination of whether a decoded message is successful relies on the use of a cyclic redundancy check (CRC) message sequence. The encoding and decoding process of the scheme with polar coding with SIC in a slot with $K$ users is demonstrated in Fig. 12.
@@ V-B4 Alternative solutions
There are also some other solutions focusing on URA over fading channels with a single-antenna receiver. For instance, in [94], the active users utilize the entire transmission time frame to transmit their data. Specifically, the time frame is divided into pilot and data parts. The active users pick a pilot sequence from a common codebook for transmission in the pilot part, and they repeat their codeword bits along the rest of the transmission time frame with permutation and scrambling in the data part. At the receiver, AMP is employed for joint pilot detection and channel estimation, and the data part is recovered by iterative data estimation and interference cancellation using the repeated codeword bits. Furthermore, the authors of [95] study URA over frequency-selective channels and propose OFDM to mitigate adverse channel effects combined with CS-based activity detection and TIN-SIC for message recovery at the receiver side.
@@
Figure 13: Performance comparison of the coding schemes in fading URA with a single-antenna receiver, with a target PUPE of $0.1$, $n=30000$, and $B=100$ (the common set of parameters in the literature).
Summary and Performance Comparison:
@@
A summary of URA studies over fading channels involving a single-antenna receiver is given in Table IV. Moreover, the performance of different coding schemes for the single-antenna fading URA setup is compared in Fig. 13 assuming $n=30000$, $B=100$, and $\epsilon=0.1$. The reason that some diagrams seem not to be fully plotted is that they diverge after a certain active user load, i.e., they cannot achieve the target PUPE no matter how much the $E_{b}/N_{0}$ is increased. As can be observed from Fig. 13, the CCS-based scheme in [89] exhibits lower energy efficiency compared to the other ones since it employs a sub-optimal coding scheme, while others use strong channel codes such as LDPC or polar codes. The performance of the LDPC-based scheme with joint decoding in [80] is inferior to that of the ones in [72], [91] and [85], which can be attributed to the absence of an efficient SIC block in this LDPC-based scheme. On the other hand, the TIN-SIC approach in [91] and [85] offers a better performance as it allows the receiver to recover the messages of users with a stronger fading coefficient, subtract their effects, and decode the remaining users in subsequent iterations, especially when it is combined with polar coding as in [85]. Nevertheless, the TIN-SIC scheme in [85] is outperformed by the spreading-based scheme in [94], and the ODMA-based scheme employing polar codes in [78], which offers the best performance among the proposed schemes, due to the sparsity of the transmissions.
Computational Complexity:
@@
We now compare the computational complexity of the single-antenna fading URA solutions described in this subsection. The tree-code-based CCS approach has a low complexity, which is linear in the list of messages at each decoding step and exponential in the number of message bits in each slot [89]. In addition, the decoding operations in [89] can be implemented by additions, which are easier to implement in hardware compared to multiplications. The complexity of the $T$-fold ALOHA-based schemes employing LDPC with joint decoding [80] or polar coding with TIN-SIC [85] is mainly determined by $T$, that is, the number of users that the decoder tries to recover in each slot. Note that TIN-SIC is more preferable than joint decoding since it utilizes single-user decoding, which makes it generally less complex and more parallelizable. Spreading-based and ODMA-based schemes, proposed in [94] and [78], respectively, have higher complexities compared to their counterparts in exchange for their superior performance. For instance, the complexity of MMSE filtering for fading coefficient estimation in [78] is cubic in $K_{a}$.
V-C MIMO Fading Channels

### v3

@@ V URA over Fading Channels
GMAC is an ideal channel model that does not capture many practical effects of the communication channel. With this motivation, more realistic fading MAC models, where the transmitted signals of the users are attenuated with random fading coefficients, are considered in several works in the URA literature.
There are some challenges in extending and applying the URA over GMAC solutions to fading scenarios. The primary challenge in fading MAC is estimating the channel coefficients at the receiver side and designing the decoder to account for fading. In the literature, two approaches are proposed to address this problem; the first is to embed the channel estimation into the decoding algorithm without requiring a separate channel estimation step (e.g., [88, 93, 115]). The second approach is to exploit a part of the frame for pilot transmission, perform explicit channel estimation using the received signal in the pilot part, and utilize the estimated coefficients for detection in the data part (e.g., [102, 120, 12]). While most schemes developed for single-antenna fading channels, and the CCS-based and tensor-based schemes in URA over MIMO fading channels opt for the former, there is also a vast literature on pilot-based schemes. The main algorithms applied for channel estimation in pilot-based schemes are MMSE [84, 120, 12, 14, 124, 119] and AMP [102, 122, 121]. Note that in most URA works over fading channels, quasi-static fading is considered.
In this section, we first review the bounds on the performance of URA solutions taking into account channel fading that have been provided in the literature. We then cover the URA solutions that present low-complexity coding schemes to achieve those bounds. To this end, in order to explain the URA approaches that take into account channel fading in a simpler setup for better understanding, we first cover the single-antenna fading solutions and then turn our attention to the MIMO fading ones.
@@ V-A Performance Bounds
In [93, 91, 92, 88], the authors derive some performance bounds for the URA framework over a Rayleigh fading channel. Specifically, the authors in [88] present a converse bound based on the results in [89] and the meta-converse bound in [90]. These results numerically demonstrate that the $E_{b}/N_{0}$ requirements for a fading channel are necessarily higher than those for the AWGN channel. They also argue that, for typical system parameters in this setup, a Fano-type converse bound is worse than the proposed one.
In [91] and [92], achievability and converse bounds are proposed for the cases of known and unknown CSI in an asymptotic setting, where the number of users grows linearly with the blocklength. The authors in [88] provide both asymptotic and non-asymptotic achievability and converse bounds for the case of unknown CSI. In [93], authors derive an achievability bound for the URA scheme over the Rayleigh fading channel by considering a slotted structure and perfect SIC in their analysis, which is tighter than the other bounds. However, since the channel coefficients used in an SIC process are estimated, hence imperfect, the assumption of a perfect SIC is not realistic, which makes the result in [93] an approximation rather than an achievability one.
Some achievability and converse bounds for URA over MIMO Rayleigh fading channels are obtained in [94, 95]. Among the developed results, the ML-based achievability bound is obtained by employing a projection decoder developed in [96]. Moreover, the results in [94] are extended to the case of random and unknown number of active users in [97].
@@ V-B Single-Antenna Fading Channels
We first examine the URA solutions over fading channels that assume a single-antenna receiver. The works in [77, 102, 104, 93, 88, 98, 99, 103, 101, 100] introduce practical coding schemes tailored for this setup. Among these works, some assume that the received signals from all users are synchronized [77, 93, 88, 98, 102, 104], while others take into account asynchronous transmission [103, 99, 101, 100]. To estimate the delay of asynchronous signals, [103, 99, 100] use OFDM-based transmission to convert the time delay to a phase shift in the frequency domain, while the authors in [101] consider time-domain transmission where the delay of a signal is detected via a correlation-based approach.
Focusing on Rayleigh fading channels, we present existing solutions for URA over fading channels, categorized based on their general approach.
@@ V-B1 CCS-based schemes
As explained in the previous section, CCS was initially proposed for URA over GMAC. In subsequent literature, it has also been applied to the fading scenario, see [98, 101, 104]. In [98], the authors replace the tree code of the GMAC CCS scheme in [63] by list recoverable codes that are capable of correcting $t$ errors. They propose two outer code constructions: The first is a modified tree code with the ability to recover $t$ errors (called the $t$-tree), and the second is a Reed-Solomon (RS)-based code. In [104], the outer encoded message bits are partitioned into sub-blocks, and each sub-block is transmitted in a different slot after CS-type encoding, with AMP at the receiver. Moreover, in [101], the authors utilize CCS for the asynchronous URA model. More specifically, during the encoding stage, they append $T_{z}$ additional zeros to the end of each CS sequence, where $T_{z}$ represents the maximum possible delay of the system. At the receiving end, the decoder takes into account all possible delays, ranging from one symbol to a maximum of $T_{z}$-symbol delays during the decoding process.
@@ V-B2 Joint decoding-based schemes
In [88], a joint decoding approach is employed. For encoding, this scheme partitions the frame into distinct slots. Each user transmits its LDPC codeword through a randomly chosen slot. As previously mentioned, slot allocation offers the significant benefit of mitigating the per-slot interference by distributing users across slots. This scheme utilizes an iterative belief propagation (BP) decoder, as illustrated in Fig. 11. The graph incorporates four distinct node types: variable nodes (in red); check nodes (in blue), which constitute the traditional LDPC decoder; functional nodes (in green) corresponding to the symbols of the received signal; and a fourth type (in magenta) corresponding to the fading coefficients. This graph allows them to update the estimated values of both the channel coefficients and the symbols simultaneously for each user.
Figure 11: Iterative BP decoder for a two-user LDPC-based scheme with joint decoding [88].
@@ V-B3 TIN-SIC-based schemes
A lower-complexity alternative to joint decoding is the TIN-based decoding. Similar to GMAC, the TIN solution for fading channels operates under the assumption that a single user is present and treats the signals of other interfering users as noise. Consequently, this approach results in an exceptionally low average SNR because it encounters interference-plus-noise as the system’s overall noise level, leading to inferior performance. However, when combined with an appropriate channel code and SIC, the performance can be improved.
@@
Paper Approach Description Synchron. vs. Asynch. Channel Estimation Channel Code
@@
Kowshik et al., 2019&2021 [91, 92] - Achievable and converse bounds for known and unknown CSI in an asymptotic setting + comparing with CDMA, TDMA, and FDMA. Synchron. - -
Kowshik et al., 2019 [100] TIN-SIC $T$-fold ALOHA-based scheme employing OFDM, channel coding, and SIC. Asynch. Embedded LDPC
Andreev et al., 2020 [93] TIN-SIC $T$-fold ALOHA-based scheme employing channel coding, single-user decoding, and SIC. Synchron. OMP Polar
@@
Kowshik et al., 2020 [88] joint decoding Achievability and converse bounds for known and unknown CSI + a scheme employing slotting, channel coding, and joint BP decoding. Synchron. Embedded LDPC
Amalladinne et al., 2019 [101] CCS Extension of CCS in [63] to accommodate asynchronous scenarios. Asynch. Embedded -
@@
Andreev et al., 2022 [98] CCS Improving CCS in [63] by replacing the outer tree code with 1) a code capable of correcting $t$ errors, and 2) a Reed–Solomon code. Synchron. Embedded RS
@@
Yun et al., 2024 [104] CCS A slotted scheme employing CCS and AMP in a block fading model. Synchron. Embedded -
@@
Nassaji et al., 2022 [102] Spreading A non-slotted scheme employing preamble transmission, channel coding, repetition, permutation, scrambling, and SIC. Synchron. AMP Polar
Ozates et al., 2024 [84] ODMA ODMA-based scheme employing gOMP for pilot and pattern detection, single-user decoding and SIC. Synchron. LMMSE Polar
@@
Ozates et al., 2023 [103] OFDM+ SIC A slotted scheme in the frequency-selective channel employing OFDM, pilot transmission, channel coding, and SIC. Asynch. OMP Polar
@@
TIN-SIC with LDPC coding: In [99], building upon the synchronous schemes presented in [88], the authors present a low-complexity coding scheme which incorporates OFDM techniques to mitigate the effects of the asynchronous transmission by converting the time delays into phase shifts, and employing SIC to decode the user packets. Further numerical results for this scheme are presented in [100].
@@
TIN-SIC with Polar Coding: In [93], a TIN-SIC solution in conjunction with polar coding is presented, where the frame in this approach is divided into $S$ slots. Each user randomly selects a slot to transmit its polar-coded and modulated signal. For decoding, the scheme considers a TIN strategy followed by SIC. Since different users in a slot experience different fading coefficients, some of them (those with larger fading coefficient amplitudes) are more likely to be decoded by TIN. After decoding the strongest user’s signal while treating the others as noise, the channel coefficient is estimated using orthogonal matching pursuit (OMP). Given the estimated channel coefficient and the decoded signal of a user, its contribution can be removed from the received signal using SIC. This strategy continues until no signal is successfully detected during an iteration. It is important to emphasize that the determination of whether a decoded message is successful relies on the use of a cyclic redundancy check (CRC) message sequence. The encoding and decoding process of the scheme with polar coding with SIC in a slot with $K$ users is illustrated in Fig. 12.
@@ V-B4 Alternative solutions
Some other solutions also focus on URA over fading channels with a single-antenna receiver. For instance, in [102], the active users utilize the entire transmission frame to transmit their data. Specifically, the frame is divided into pilot and data parts. The active users select a pilot sequence from a common codebook for transmission in the pilot part, and they repeat their codeword bits for the remainder of the transmission time frame, with permutation and scrambling applied in the data part. At the receiver, AMP is employed for joint pilot detection and channel estimation, and the data part is recovered by iterative data estimation and interference cancellation using the repeated codeword bits.
The ODMA approach is also applied to the single-antenna fading scenario in [84] by allocating a portion of the frame for pilot transmission while a polar codeword is transmitted in the rest of the frame by ODMA. At the receiver side, the selected pilots and patterns are detected by a generalized OMP (gOMP) algorithm, followed by linear MMSE (LMMSE) channel estimation, single-user decoding, and SIC to recover the transmitted messages. Furthermore, the authors in [103] study URA over frequency-selective channels and propose OFDM to mitigate adverse channel effects combined with CS-based activity detection and TIN-SIC for message recovery at the receiver side.
@@
Figure 13: Performance comparison of the coding schemes for URA over fading channel with a single-antenna receiver, with a target PUPE of $0.1$, $n=30000$, and $B=100$ (the common set of parameters in the literature).
V-C Summary and Comparison of Single-Antenna Fading Solutions
A summary of URA studies over fading channels involving a single-antenna receiver is given in Table IV, where Embedded under channel estimation column refers to the case that estimation of the fading coefficient is included as a part of the decoding algorithm.
V-C1 Performance Comparison:
@@
The performance of different coding schemes developed for URA over single-antenna fading channels is compared in Fig. 13 assuming $n=30000$, $B=100$, and $\epsilon=0.1$. The reason that some diagrams seem not to be fully plotted is that those solutions do not work after a certain active user load, i.e., they cannot achieve the target PUPE, no matter how much the $E_{b}/N_{0}$ is increased. As can be observed from Fig. 13, the CCS-based scheme in [98] exhibits lower energy efficiency compared to the other alternatives, as it employs a suboptimal coding scheme, whereas the others use strong channel codes, such as LDPC or polar codes. The performance of the LDPC-based scheme with joint decoding in [88] is inferior to that of the ones in [77], [100], and [93], which can be attributed to the absence of an efficient SIC block in this LDPC-based approach. On the other hand, the TIN-SIC approach in [100] and [93] offers a better performance as it allows the receiver to recover the messages of users with a stronger fading coefficient, subtract their effects, and decode the remaining users in subsequent iterations, especially when combined with polar coding as in [93]. We also note that the TIN-SIC scheme in [93] is outperformed by the spreading-based solution in [102], and the ODMA-based scheme employing polar codes in [84], which offers the best performance among the proposed schemes, due to the sparsity of the transmissions.
Overall, the results in fading channels show that CCS-based solutions lack a strong channel code to provide a high energy efficiency. Simple TIN-SIC solutions offer a good performance in low user loads, however, they demonstrate the limitations of weak interference management as device density grows. The comparative analysis suggests that practical single-antenna fading URA systems will favor ODMA and spreading-based approaches, which can be flexibly adapted to varying user loads without incurring excessive complexity.
V-C2 Computational Complexity Comparison
@@
We now compare the computational complexity of the URA solutions over single-antenna fading channels. The tree-code-based CCS approach has a low complexity, which is linear in the size of the list of messages at each decoding step and exponential in the number of message bits in each slot [98]. In addition, the decoding operations in [98] can be carried out by additions, which are easier to implement on hardware compared to multiplications. The complexity of the $T$-fold ALOHA-based schemes employing LDPC with joint decoding [88] or polar coding with TIN-SIC [93] is mainly determined by the parameter $T$, that is, the number of users that the decoder tries to recover in each slot. Note that TIN-SIC is more preferable than joint decoding since it utilizes single-user decoding, which makes it generally less complex and more parallelizable. Spreading-based and ODMA-based schemes, proposed in [102] and [84], respectively, have higher complexities compared to their counterparts in exchange for their superior performance. For instance, the complexity of MMSE filtering for fading coefficient estimation in [84] is cubic in $K_{a}$.
V-D MIMO Fading Channels